### v1

@@ 6.1.1.4   Cosmic void
The next-generation survey telescopes such as CSST will probe more than ten thousand square degrees of sky area with great depth and sensitivity. Hundreds of thousands of cosmic voids are expected to be identified in the CSST spectroscopic survey at $1.50<z<$ 1.5. Using this spec-$z$ void sample, we can precisely measure the void size functions [45], void number counts [46], and void power spectra [47] at different redshifts for investigating the equation of state of dark energy and other cosmological parameters.
@@ 6.1.1.5   SN Ia
Using simulated CSST-SC SN Ia observational data, it is found that after removing the contamination of core-collapse supernovae (CCSNe) through methods such as light curve fitting, the photometric observations of the CSST UDF can obtain high-quality data of light curves of more than 2,000 SNe Ia within two years. It is found that the fraction of high-$z$ SNe Ia in the CSST UDF survey is $\sim 80\%$ at $z>0.5$ and $\sim 15\%$ at $z>1$, which is is very advantageous in constraining the dynamical evolution of dark energy. Using this SN Ia sample, the precision of constraints on cosmological parameters such as the equation of state of dark energy can be improved by almost twice compared to the current SN Ia observation data [50].
In addition, the CSST XDF planned by CSST-MCI also will make significant progress in SN Ia cosmology. The CSST-MCI can capture full-band light variation events with a time scale of more than 5 minutes (1 hour) for astronomical objects at a depth of at least 26 (27) magnitude, which is very helpful in capturing high-$z$ supernovae in the CSST XDF, and can quickly and simultaneously determine their coordinates and approximate redshift information. The CSST-MCI XDF observation is expected to discover 100-200 SNe Ia at $1.51.0<z<$ 1.5 in one year (2 hours of deep exposure per week), which can help us gain a decisive understanding of the equation of state of dark energy.
@@ 6.1.1.6   BAO and RSD
BAO is caused by the fact that in the early Universe, baryons and photons were coupled together, and under gravity and pressure, an oscillation similar to a sound wave was formed. This oscillation caused the local matter fluctuation formed and propagated outward like a sound wave, i.e. BAO. It can be seen as the “standard ruler” for measuring the cosmological distance, which is roughly 150 Mpc, and its size depends on cosmological parameters, especially the equation of state of dark energy.
@@
The CSST slitless spectroscopic survey can observe 17,500 deg2 sky area, and obtain more than one hundred million galaxy spectra and about four million AGN spectra [52]. This indicates that CSST can not only measure BAO with galaxies but also with AGNs over a larger redshift range. By considering the reconstruction technique, it is found that the CSST spec-$z$ survey can measure the two BAO scaling parameters $\alpha_{\parallel}$ and $\alpha_{\perp}$ with precisions $\sim 1\%$ using galaxies at $1.20.3<z<$ 1.2. When using AGNs, the precisions can reach $\sim 3\%$ at $z<3$. By combining the BAO data from CSST spec-$z$ galaxy and AGN surveys, the constraint accuracy of $w$ is expected to achieve $\sim 9\%$ [52].
@@
For the RSD observation, based on theoretical estimation, CSST can measure the structure formation factor $f\sigma_{8}$ with a precision of $\sim 2\%$ at $10.4<z<$ 1, where $f$ is the growth rate. Compared with DESI, CSST RSD observation can constrain $f\sigma_{8}$ much more accurately at $z<0.6$. This means that the CSST spec-$z$ survey can be complementary to other spec-$z$ surveys such as DESI. By combining the CSST and other spec-$z$ surveys, we expect to discover evidence of the evolution of dark energy at a high level of confidence.
@@ 6.1.1.7   Strong gravitational lensing
The strong gravitational lensing systems detected by CSST also can be used to constrain the property of dark energy [56]. The most commonly used strong lensing methods include the velocity dispersion of lens galaxy, the time delay between multiple images, etc.
@@
It is expected that, about 60% of the strong lensing systems observed by CSST are located at $z>0.5$, and about 15% at $21<z<$ 2 [60]. Therefore, The MSP system in the CSST strong lensing sample can provide a supplementary tool for studying the proportion and evolution of dark matter components in the Universe at $z<2$. Besides, we can not only use MSP systems independently, but also combine the MSP system with other cosmological probes (such as CMB, BAO, time delay, etc.) to more accurately constrain $\Omega_{\rm m}$ and $w$.
@@ 6.1.2.2   Axion dark matter
Leaving aside the debate over the “temperature” of dark matter, the CDM model itself also contains a variety of different candidates, such as the axion. The axion is a hypothetical scalar particle proposed in the 1970s to solve the strong CP problem in quantum chromodynamics. The mass of the axion is weakly constrained in theory, and has a very wide possible mass range. The axion is a good candidate particle for cold dark matter.
In cosmology, the ultra-light axion (ULA) with a mass of about $10^{-22}$ eV has attracted much attention. For ULA dark matter, due to the extremely small particle mass, its de Broglie wavelength can reach the order of magnitude of the dwarf galaxy or galaxy scale, so the ULA is consistent with the prediction of the standard CDM model at large scales ($>10$ kpc); but at smaller scales (less than the Jeans length), the ULA dark matter model shows some unique characteristics.
@@ 6.1.2.3   Self-interacting dark matter
For a fundamental particle, the two most important properties are mass and interaction cross section. It is known that dark matter basically does not interact with visible matter, so do dark matter particles collide with themselves, and what is the cross section? These questions are crucial for understanding the nature of dark matter. Self-interacting dark matter (SIDM) has a self-scattering cross section that is comparable to the nucleon-nucleon scattering cross section.
@@
In astronomy, merging galaxy clusters are natural large particle colliders, providing ideal laboratories for measuring dark matter self-interactions. During the process of galaxy cluster merging, the non-gravitational interactions between dark matter and standard model particles will be more significant. In addition to galaxy cluster mergers, we can also study SIDM at the scale of galaxy clusters and their substructures. Considering that cluster subhalos and main cluster halos probe different regions of velocity space, these studies at different scales are helpful for exploring the potential velocity dependence of the self-interaction cross section of dark matter. These effects can be precisely measured in the CSST gravitational lensing and galaxy cluster observations.
@@ 6.1.3 Neutrino
Using the CSST 3$\times$2pt data, we can effectively constrain $\Sigma m_{\nu}$, considering other cosmological and systematical parameters. In Figure 7, the constraint result of $\Sigma m_{\nu}$ and some of other cosmological parameters have been shown. The systematics from shear and photo-$z$ calibration, intrinsic alignment, galaxy bias, and baryonic feedback effect have been considered in the analysis. It is found that the CSST 3$\times$2pt surveys can provide a stringent constraint on the total neutrino mass with $\Sigma m_{\nu}\lesssim 0.36\ (0.56)$ eV at 68% (95 %) CL. [36]. By including other cosmological data, such as CMB, CSST can measure $\Sigma m_{\nu}$ with extremely high precision, and can decisively distinguish the neutrino mass hierarchy.
@@ 6.1.4 Gravity
A series of experiments in the inner Solar System and observations of neutron stars have shown that Einstein’s general relativity is strictly applicable in these two environments. Therefore, any modified gravity theory must rely on a special “screening mechanism” to fully restore general relativity in areas of high density or high gravitational potential. The currently popular screening mechanisms of modified gravity can be roughly divided into two categories: the first is the thin-shell screening mechanism based on the strength of the scalar field, such as the chameleon mechanism, the symmetrone mechanism, and the dilaton mechanism; the second category is the kinetic screening mechanism based on the first or second derivatives of the scalar field, such as the Vainshtein mechanism and the K-mouflage mechanism. These two types of shielding mechanisms will have different predictions for the redshift and scale dependence of cosmic structure growth. By combining cosmological numerical simulations based on different modified gravity theories, the CSST survey can provide high-precision constraints on the modified gravity models.
The popular modified gravity theories in cosmological research include the $f(R)$ and $f(T)$ models based on the chameleon mechanism, and the DGP, Horndeski and Galileon models based on the Vainshtein mechanism, etc. For example, the $f(R)$ theory is a generalization of the Einstein-Hilbert action, which generalizes the action from the Ricci scalar to an arbitrary function of the Ricci scalar. For the translational gauge theory of gravity equivalent to general relativity, there is also a similar $f(T)$ theory that generalizes the torsion scalar to an arbitrary function of the torsion scalar [68].
The CSST weak gravitational lensing or shear measurement can directly probe the dependence of dark matter clustering on scale and redshift in the Universe. The matter clustering predicted by most modified gravity theories is on a near-linear scale ($0.2-20\ h^{-1}$pc), and the S/N of the shear signal is strongest at this scale. Therefore, if the CSST shear measurement has any statistically significant deviation from the predictions of the general relativity+dark energy model, we believe that there may be evidence of modified gravity on the relevant scale. However, the baryonic process and neutrinos may also have a large impact on the matter power spectrum at this scale, thereby weakening the constraints of cosmic shear on modified gravity. In addition, we can also use the stronger signal in cosmic shear, i.e. the peak statistics of weak gravitational lensing, to constrain gravity models. For the $f(R)$ theory, it is found that the CSST weak gravitational lensing measurement can constrain $|f_{R0}|=|{\rm d}f/{\rm d}R|_{z=0}$ to an accuracy of $\sim 10^{-6}$. Besides, using the CSST SN Ia and BAO data, we can also obtain stringent constraint on the $f(R)$ theory [69].
@@
Additionally, galaxy clusters also can be used to test the gravity theory and study the modified gravity. CSST wide-field multi-band photometric survey can detect approximately 300,000 galaxy clusters with a mass greater than $10^{14}\ M_{\odot}$ at $z<1.5$, providing an excellent sample of massive dark matter halos for measuring the structure growth history of the Universe and testing modified gravity theories. Using the shear signal of galaxy clusters, CSST will accurately measure the density profiles of galaxy clusters with different satellite galaxy abundances between 0.1 $h^{-1}$pc and 30 $h^{-1}$pc, thereby detecting the density enhancement effect caused by modified gravity near the Virial radius of galaxy clusters [72]. The CSST weak gravitational lensing survey can also measure the “splashback radius” of the dark halos of galaxy clusters with an accuracy of one percent, which can be used to study the modified gravity theories [73]. Besides, combined with data from spectroscopic surveys, CSST can also use the Galaxy Infall Kinematics (GIK) method to test modified gravitational theories [74].
@@ 6.1.5 Primordial perturbations
The primordial metric perturbations of the Universe contain information about the early Universe, and are of great significance for exploring the origin of the Universe and the physical laws at ultra-high energy scales. In the standard model, the large-scale primordial metric perturbations of the Universe originate from the inflation. The power spectrum of the primordial perturbations given by the simplest single-field slow-roll inflation, referred to as the “primordial power spectrum”, is very close to a power-law, which is also supported by the CMB observations.
@@
Compared with the two-dimensional CMB data, the matter density field of the Universe is three-dimensional, and in principle contains more information, especially at linear scales. For instance, the primordial non-Gaussianity can affect the galaxy bias and make it scale-dependent at linear scales [75], which can be used to constrain $f_{\rm NL}^{\rm local}$. Therefore, at least at linear scales, it is reliable to use the LSS or galaxy clustering data to constrain the property of primordial perturbations. Using the galaxy clustering data in the CSST spectroscopic survey, it is found that the $1\sigma$ error of $f_{\rm NL}^{\rm local}$ can reach $\sim 6$, which is comparable to the CMB result [76].
The scalar spectral index $n_{\rm s}$ and the running of the spectral index $\alpha_{\rm s}={\rm d}n_{\rm s}/{\rm d\,ln}k$ of the primordial power spectrum also can be tightly constrained by the CSST 3$\times$2pt, peak statistics of weak gravitational lensing, and galaxy spec-$z$ clustering observations, with an accuracy higher than 3% [15, 36, 35, 37].
@@ 6.1.6 Hubble parameter
In addition to the probes mentioned above, other cosmological probes, such as AP effect [83], galaxy-ellipticity correlation [84], and cosmic optical background [85], also can be used in the CSST photometric and spectroscopic surveys to test cosmological models or theories, and constrain the cosmological parameters. CSST also has excellent synergy with other telescopes focusing on different bands for the cosmological study, such as CMB experiments [86], radio telescopes like Five-hundred-meter Aperture Spherical Telescope (FAST) [87] and MeerKAT [88] for detecting 21-cm signal, and the future gravitational wave detectors [89]. Considering all these observations, CSST is expected to make great contributions in the above cosmological fields.
@@ 6.2.1 AGN and massive black hole
Figure 8: The QSO spectra at $z=0.8$ (blue dashed curve) and $z=2.2$ (red solid curve), and F-type star spectrum (gray solid curve), compared to the CSST transmission curves of the seven photometric bands (i.e. $NUV$, $u$, $g$, $r$, $i$, $z$, and $y$). The $NUV$ and $u$ bands of the CSST are important for estimating photo-$z$ of quasars and distinguishing them from stars.
@@
Third, with the help of multi-band data provided by CSST and other telescopes, special AGNs and quasars (such as changing-look AGNs) can be searched and confirmed. Their physical origins and evolution history will be explored, and new observational support for black hole growth, galaxy feedback mechanisms, and large-scale structure formation can be provided. Using the CSST multi-color photometric and slitless spectroscopic data, as well as ground-based and space-based multi-band data (such as $Euclid$, RST etc.), CSST will discover a large number of high-$z$ quasars ($\sim 4000$ at $z\simeq 6$ and $\sim 700$ at $z\simeq 7$) with the blue-end dropout technology (see Figure 9). The CSST wide-field survey will detect tens of thousands of lensed quasar candidates and further confirm strong lensed quasars through follow-up observations to explore a series of scientific issues involving foreground galaxies, quasar physical properties, and Hubble constant measurements. CSST also can detect thousands of new changing-look AGNs and a batch of narrow-line Seyfert 1 galaxies (NLS1s) and blazers, and will greatly advance the studies of these special AGNs.

### v3

@@ 6.1.1.4 Cosmic void
The upcoming Stage-IV survey telescopes such as CSST will probe more than ten thousand square degrees of sky area with great depth and sensitivity. Hundreds of thousands of cosmic voids are expected to be identified in the CSST spectroscopic survey at $0<z<1.5$ [58, 59]. Using this spec-$z$ void sample, we can precisely measure the void size functions [58], void number counts [59], and void power spectra [60] at different redshifts for investigating the equation of state of dark energy and other cosmological parameters.
@@ 6.1.1.5 SN Ia
Using simulated CSST-SC SN Ia observational data, it is found that after removing the contamination of core-collapse supernovae (CCSNe) through methods such as light curve fitting, the photometric observations of the CSST UDF can obtain high-quality data of light curves of more than 2,000 SNe Ia within two years [63]. It is found that the fraction of high-$z$ SNe Ia in the CSST UDF survey is $\sim 80\%$ at $z>0.5$ and $\sim 15\%$ at $z>1$, which is is very advantageous in constraining the dynamical evolution of dark energy. Using this SN Ia sample, the precision of constraints on cosmological parameters such as the equation of state of dark energy can be improved by almost twice compared to the current SN Ia observation data [63].
In addition, the CSST XDF planned by CSST-MCI also will make significant progress in SN Ia cosmology. The CSST-MCI can capture full-band light variation events with a time scale of more than 5 minutes (1 hour) for astronomical objects at a depth of at least 26 (27) magnitude, which is very helpful in capturing high-$z$ supernovae in the CSST XDF, and can quickly and simultaneously determine their coordinates and approximate redshift information. The CSST-MCI XDF observation is expected to discover 100-200 SNe Ia at $1.0<z<1.5$ in one year (2 hours of deep exposure per week), which can help us gain a decisive understanding of the equation of state of dark energy.
@@ 6.1.1.6 BAO and RSD
BAO is caused by the fact that in the early Universe, baryons and photons were coupled together, and under gravity and pressure, oscillation was formed. This oscillation caused the local matter fluctuations formed and propagated outward as sound waves, i.e. BAO. It can be seen as the “standard ruler”, which is roughly 150 Mpc, for measuring the cosmological distance and the property of dark energy.
@@
The CSST slitless spectroscopic survey can observe 17,500 deg2 sky area, and obtain more than one hundred million galaxy spectra and about four million AGN spectra [65]. This indicates that CSST can not only measure BAO with galaxies but also with AGNs over a larger redshift range. By considering the reconstruction technique, it is found that the CSST spec-$z$ survey can measure the two BAO scaling parameters $\alpha_{\parallel}$ and $\alpha_{\perp}$ with precisions $\sim 1\%$ using galaxies at $0.3<z<1.2$. When using AGNs, the precisions can reach $\sim 3\%$ at $z<3$. By combining the BAO data from CSST spec-$z$ galaxy and AGN surveys, the constraint accuracy of $w$ is expected to achieve $\sim 9\%$ [65].
@@
For the RSD observation, based on theoretical estimation, CSST can measure the structure formation factor $f\sigma_{8}$ with a precision of $\sim 2\%$ at $0.4<z<1$, where $f$ is the growth rate. Compared with DESI, CSST RSD observation can constrain $f\sigma_{8}$ more accurately by a factor of 2 at least at $z<0.6$ [69], since CSST spectroscopic survey has larger survey area and deeper magnitude limit. This means that the CSST spec-$z$ survey can be complementary to other spec-$z$ surveys such as DESI. By combining the CSST and other spec-$z$ surveys, we expect to measure the equation of state of dark energy at a high level of precision in a large redshift range.
@@ 6.1.1.7 Strong gravitational lensing
The strong gravitational lensing systems detected by CSST also can be used to constrain the property of dark energy [70]. There are several commonly used strong lensing methods in cosmological study, such as the velocity dispersion of lens galaxies and time delay between multiple images.
@@
It is expected that, about 60% of the strong lensing systems observed by CSST are located at $z>0.5$, and about 15% at $1<z<2$ [74]. Therefore, The MSP system in the CSST strong lensing sample can provide a supplementary tool for studying the proportion and evolution of dark matter components in the Universe at $z<2$. Besides, we can not only use MSP systems independently, but also combine the MSP system with other cosmological probes (such as CMB, BAO, time delay, etc.) to more accurately constrain $\Omega_{\rm m}$ and $w$.
@@ 6.1.2.2 Axion dark matter
Leaving aside the debate over the “temperature” of dark matter, the CDM model itself also contains a variety of different candidates, such as the axion. The axion is a hypothetical scalar particle proposed in the 1970s to solve the strong CP problem in quantum chromodynamics [76, 77, 78]. The mass of the axion is weakly constrained in theory, and has a very wide possible mass range. The axion is a good candidate particle for cold dark matter.
In cosmology, the ultra-light axion (ULA) with a mass of about $10^{-22}$ eV has attracted much attention. For ULA dark matter, due to the extremely small particle mass, its de Broglie wavelength can reach the order of magnitude of the dwarf galaxy or galaxy scale, so the ULA is consistent with the prediction of the standard CDM model at large scales (>10>10 kpc); but at smaller scales, the ULA dark matter model shows some unique characteristics. It can affect the transfer function of matter power spectrum and suppress the small-scale structure formation.
@@ 6.1.2.3 Self-interacting dark matter
For a fundamental particle, the two most important properties are mass and interaction cross section. It is known that dark matter basically does not interact with visible matter, so do dark matter particles collide with themselves, and what is the cross section? These questions are crucial for understanding the nature of dark matter. Self-interacting dark matter (SIDM) has a self-scattering cross section that is comparable to the nucleon-nucleon scattering cross section [79].
@@
In astronomy, merging galaxy clusters are natural large particle colliders, providing ideal laboratories for measuring dark matter self-interactions [80, 81]. During the process of galaxy cluster merging, the non-gravitational interactions between dark matter and standard model particles will be more significant. In addition to galaxy cluster mergers, we can also study SIDM at the scale of galaxy clusters and their substructures. Considering that cluster subhalos and main cluster halos probe different regions of velocity space, these studies at different scales are helpful for exploring the potential velocity dependence of the self-interaction cross section of dark matter. These effects can be precisely measured in the CSST gravitational lensing and galaxy cluster observations.
@@ 6.1.3 Neutrino
[…]
@@ 6.1.4 Gravity
A series of experiments in the inner Solar System and observations of binary pulsar timing have shown that Einstein’s general relativity is strictly applicable in these two environments. Therefore, any modified gravity theory must rely on a special “screening mechanism” to fully restore general relativity in areas of high density or high gravitational potential. The currently popular screening mechanisms of modified gravity can be roughly divided into two categories: the first is the thin-shell screening mechanism based on the strength of the scalar field, such as the chameleon mechanism, the symmetron mechanism, and the dilaton mechanism; the second category is the kinetic screening mechanism based on the first or second derivatives of the scalar field, such as the Vainshtein mechanism and the K-mouflage mechanism. These two types of shielding mechanisms will have different predictions for the redshift and scale dependence of cosmic structure growth. By combining cosmological numerical simulations based on different modified gravity theories, the CSST survey can provide high-precision constraints on the modified gravity models.
The popular modified gravity theories in cosmological research include the models of $f(R)$, $f(T)$, DGP, Galileon, Horndeski, etc. For example, the $f(R)$ theory is a generalization of the Einstein-Hilbert action, which generalizes the action from the Ricci scalar to an arbitrary function of the Ricci scalar. For the translational gauge theory of gravity equivalent to general relativity, there is also a similar $f(T)$ theory that generalizes the torsion scalar to an arbitrary function of the torsion scalar [91].
The CSST weak gravitational lensing or shear measurement can directly probe the dependence of dark matter clustering on scale and redshift in the Universe. The matter clustering predicted by most modified gravity theories is on a near-linear scale, and the S/N of the shear signal is strongest at this scale. Therefore, if the CSST shear measurement has any statistically significant deviation from the predictions of the general relativity+dark energy model, we believe that there may be evidence of modified gravity on the relevant scale. However, the baryonic process and neutrinos may also have a large impact on the matter power spectrum at this scale, thereby weakening the constraints of cosmic shear on modified gravity. In addition, we can also use the stronger signal in cosmic shear, i.e. the peak statistics of weak gravitational lensing, to constrain gravity models. For the $f(R)$ theory, it is found that the CSST weak gravitational lensing measurement can constrain $|f_{R0}|=|{\rm d}f/{\rm d}R|_{z=0}$ to an accuracy of $\sim 10^{-6}$. Besides, using the CSST SN Ia and BAO data, we can also obtain stringent constraint on the $f(R)$ theory [92].
@@
Additionally, galaxy clusters also can be used to test the gravity theory and study the modified gravity. CSST wide-field multi-band photometric survey can detect approximately 300,000 galaxy clusters with a mass greater than $10^{14}\ M_{\odot}$ at $z<1.5$, providing an excellent sample of massive dark matter halos for measuring the structure growth history of the Universe and testing modified gravity theories. Using the shear signal of galaxy clusters, CSST will accurately measure the density profiles of galaxy clusters with different satellite galaxy abundances between 0.1 $h^{-1}$Mpc and 30 $h^{-1}$Mpc, thereby detecting the density enhancement effect caused by some modified gravity models near the Virial radius of galaxy clusters [95]. The CSST weak gravitational lensing survey can also measure the “splashback radius” of the dark halos of galaxy clusters with an accuracy of one percent, which can be used to study the modified gravity theories [96]. Besides, combined with data from spectroscopic surveys, CSST can also use the Galaxy Infall Kinematics (GIK) method to test modified gravitational theories [97].
@@ 6.1.5 Primordial perturbations
The primordial metric perturbations of the Universe contain information about the early Universe, and are of great significance for exploring the origin of the Universe and the physical laws at ultra-high energy scales. In the standard model, the large-scale primordial metric perturbations of the Universe originate from the inflation [98, 99]. The power spectrum of the primordial perturbations given by the simplest single-field slow-roll inflation, referred to as the “primordial power spectrum”, is very close to a power-law, which is also supported by the CMB observations.
@@
Compared with the two-dimensional CMB data, the matter density field of the Universe is three-dimensional, and in principle contains more information, especially at linear scales. For instance, the primordial non-Gaussianity can affect the galaxy bias and make it scale-dependent at linear scales [100], which can be used to constrain $f_{\rm NL}^{\rm local}$. Therefore, at least at linear scales, it is reliable to use the LSS or galaxy clustering data to constrain the property of primordial perturbations. Using the galaxy clustering data in the CSST spectroscopic survey [15], it is found that the $1\sigma$ error of $f_{\rm NL}^{\rm local}$ can reach $\sim 6$, which is comparable to the CMB result [101].
[…]
@@ 6.1.6 Hubble parameter
In addition to the probes mentioned above, other cosmological probes, such as AP effect [108], galaxy-ellipticity correlation [218], and cosmic optical background [110], also can be used in the CSST photometric and spectroscopic surveys to test cosmological models or theories, and constrain the cosmological parameters.
CSST also has excellent synergy with other telescopes in the cosmological study. Compared to other Stage-IV survey telescopes, such as $Euclid$, RST and LSST, CSST has some significant advantages. For example, CSST-SC has high dynamic image quality ($R_{\rm EE80}<0.15^{\prime\prime}$), near-UV observation ($NUV$ band), and multi-band coverage (7 photometric and 3 spectroscopic bands), which can obtain high-resolution images, accurate photo-$z$ estimation, and adequate spectral information for cosmological and other studies. This provides a good foundation for cooperation with other Stage-IV telescopes, and has excellent complementarity in observational bands (e.g. near-UV and near-IR bands) and other aspects [14, 45]. In addition, CSST can also conduct joint observations with telescopes or detectors in other bands or methods, such as CMB experiments [111], radio telescopes like Five-hundred-meter Aperture Spherical Telescope (FAST) [112] and MeerKAT [113] for detecting 21-cm signal, and the future gravitational wave detectors [114]. Considering all these observations, CSST is expected to make great contributions in the above cosmological fields.
@@ 6.2.1 AGN and massive black hole
Figure 8: The QSO spectra at $z=0.8$ (blue dashed curve) and $z=2.2$ (red solid curve), and F-type star spectrum (gray solid curve), compared to the CSST transmission curves of the seven photometric bands (i.e. $NUV$, $u$, $g$, $r$, $i$, $z$, and $y$). The NUVNUV and $u$ bands of the CSST are important for estimating photo-$z$ of quasars and distinguishing them from stars.
@@
Third, with the help of multi-band data provided by CSST and other telescopes, special AGNs and quasars (such as changing-look AGNs) can be searched and confirmed. Their physical origins and evolution history will be explored, and new observational support for black hole growth, galaxy feedback mechanisms, and large-scale structure formation can be provided. Using the CSST multi-color photometric and slitless spectroscopic data, as well as ground-based and space-based multi-band data (such as $Euclid$, RST etc.), CSST will discover a large number of high-$z$ quasars ($\sim 4000$ at $z\simeq 6$ and $\sim 700$ at $z\simeq 7$) with the blue-end dropout technology (see Figure 9), based on the luminosity functions of quasars at $z\simeq 6$ and $z>6.5$ [115, 116]. The CSST wide-field survey will detect tens of thousands of lensed quasar candidates and further confirm strong lensed quasars through follow-up observations to explore a series of scientific issues involving foreground galaxies, quasar physical properties, and Hubble constant measurements. CSST also can detect thousands of new changing-look AGNs and a batch of narrow-line Seyfert 1 galaxies (NLS1s) and blazars, and will greatly advance the studies of these special AGNs.